3 d^ 2 y dx^ 2 + dy dx -14y=0;y(0)=1,y^ prime (0)= -1

1
Expert's answer
2021-12-21T12:22:30-0500

"3y''+y'-14=0\\
\textsf{The auxiliary equation is}\\
3m²+m-14\\
m_1=2, m_2=\frac{-7}{3}\\
\textsf{Hence, the general solution is}\\
y=Ae^{2x}+Be^{\frac{-7}{3}x}\\
\therefore y'=2Ae^{2x}-\frac73Be^{\frac{-7}{3}x}\\
\textsf{Given,}\\

y(0)=1, y'(0)=-1\\
\textsf{substituting the given conditions, we have}\\
1=A+B\\
B=1-A\\
\textsf{Also,}\\
-1=2A-\frac73B\\
-3=6A-7+7A\\
13A=4\\
A=\frac4{13},\\

B=\frac9{13}\\
\textsf{Hence, the particular solution is}\\
y=\frac4{13}e^{2x}+\frac9{13}e^{-\frac73x}\\"